Jenny's utility function is given below. Jenny has $12 from her job to spend. Good 1 cost some value greater than $0, good 2 cost $1, and good 3 cost $3

u(x1, x2, x3)= min(x1, x2)+ x3

a) What is Jenny's demand for good 1, x1*(p1)

b) What is Jenny's demand curve for good 1.
Answer #1

Q B = $1 and P= $3 $12, Income = Solut * © U(21X2 ,lb= min (201 X2 ) fXj Assumine In perfect complements b/w X and X2, 24 = xX, * (b 12-3x3 trual (if pu) Supposo, Po=2 then x²= 12-312 12 * X 4 - Xa | for p 2 The dem and are for good I will be straigh